OpenStudy (anonymous):

limx-0 [x^2/sinxtanx] where [.] denotes greatest integer function answer is zero but i want the method to solve this problem

OpenStudy (anonymous):

maybe L'hopital's I tried it without breaking up tanx into sinx/cosx but that didn't work maybe try breaking it up to get x^2/(sin^2x/cosx)

OpenStudy (anonymous):

looking easy but it is difficult

OpenStudy (anonymous):

\[ \frac {x^2} {\sin x \,\tan x}= \frac{x} {\sin x} \frac{x} {\tan x} \] So the limit is 1 since each factor of the product above goes to 1 when x goes to zero.